OpenStudy (drosetelaya4life):

If p(x) = 2x2 – 4x and q(x) = x – 3, what is (p*q)(x)

OpenStudy (jdoe0001):

(p*q)(x) simply means p(x) * g(x) which in this case will be \(\bf (2x^2-4x) \cdot (x-3)\)

OpenStudy (sunnnystrong):

Recall that: \[(p*q)(x)=p(x)*q(x)\] if \[p(x)=2x^2-4x\] \[g(x)=x-3\] than: \[[2x^2-4x][x-3]=2x^3-2x^2-4x^2+12x=2x^3-6x^2+12x\] \[(p*q)(x)=2x^3-6x^2+12x\]